# SenseShift6D: Multimodal RGB-D Benchmarking for Robust 6D Pose Estimation across Environment and Sensor Variations

### v1

@@ 4 Experiments
To assess how sensor control improves efficiency in 6D pose estimation relative to conventional training and inference approaches, we conducted experiments on several state-of-the-art models.
4.1 Experiment Setup
Models. We chose three 6D object pose estimation models: ZebraPose [22], GDRNPP [19], and HiPose [17]. ZebraPose is an RGB-based method that leverages features extracted from a ResNet-34 backbone. GDRNPP improves upon its predecessor, GDR-Net  [25], by adopting a ConvNeXt  [20] backbone and modifying training details such as learning rate scheduling and weight decay. While GDRNPP includes a depth-based pose refinement module, its inference primarily relies on RGB inputs. HiPose is an RGB-D model that uses ConvNeXt as the RGB backbone and RandLA-Net  [13] to extract point cloud features from depth. These RGB and depth features are fused through a bidirectional interaction module for pose estimation. In our experiments, we excluded the refinement stage for ZebraPose and GDRNPP, while evaluating the full pipeline of HiPose.
@@ 4.1 Experiment Setup
Train setup. For training, we first generated 10k photorealistic (PBR) synthetic images per object using BlenderProc [4]. Each model was trained independently on a per-object basis, following one of the three schemes: (1) PBR + Train-Def (default physical images only), (2) PBR + Train-Def with data augmentation (including Gaussian blur, linear contrast adjustment, brightness enhancement and others – see Appendix A for details), and (3) PBR + Train-Var (full sensor and lighting variation). ZebraPose and GDRNPP were trained using an NVIDIA RTX 3090 GPU, while HiPose was trained on an NVIDIA A100 GPU. During training, we utilized depth images captured under the default-capture mode.
Metric. We used the standard Average Distance (ADD) metric [8] to evaluate model performance. ADD measures the average point-wise distance between model points transformed by the predicted pose and the ground-truth pose:
$\text{ADD}=\frac{1}{m}\sum_{\mathbf{x}\in\mathcal{M}}\left\|(\mathbf{R}\mathbf{x}+\mathbf{T})-(\tilde{\mathbf{R}}\mathbf{x}+\tilde{\mathbf{T}})\right\|,$ (1)
where $\mathcal{M}$ denotes the set of 3D model points, $\mathbf{x}\in\mathcal{M}$ represents an individual point on the model, $m$ is the total number of points in the model. $\mathbf{R}$ and $\mathbf{T}$ are the ground-truth rotation matrix and translation vector, while $\tilde{\mathbf{R}}$ and $\tilde{\mathbf{T}}$ are their predicted counterparts.
4.2 Analysis on RGB Sensor Control
Table 4 evaluates the impact of illumination and RGB sensor variations on 6D pose estimation. All reported results utilize the ADD metric, using AR@5, which corresponds to recall at a threshold of 5% of the object diameter. We compare three sensor control settings: (1) AE (auto-exposure) represents the conventional baseline with automatic sensor settings; (2) Rand (random exposure and gain) indicates average performance across randomly selected sensor configurations; (3) Oracle represents the best achievable performance through optimal sensor control – measuring the probability that at least one sensor-controlled configuration enables correct pose prediction for a given scene.
@@ 4.2 Analysis on RGB Sensor Control
Table 4: Performance of various sensor control methods under different training and test settings of SenseShift6D.
ZebraPose [22] GDRNPP [19] HiPose [17]
@@
Object Train AE AE Rand Oracle AE AE Rand Oracle AE AE Rand Oracle
Spray PBR + Train-Def 100.0 96.09 52.26 99.51 100.0 96.59 52.02 99.02 100.0 98.73 60.00 100.0
@@
PBR + Train-Var 90.47 89.04 68.35 95.23 90.48 90.004 72.22 97.14 97.62 97.62 86.86 98.10
Tincase PBR + Train-Def 95.55 84.43 44.15 96.00 82.22 75.11 43.29 90.67 73.33 60.36 40.27 77.78
PBR + Train-Def w/ Aug 95.55 88.88 53.37 99.10 73.33 69.78 48.85 94.67 75.56 72.44 51.09 76.00
PBR + Train-Var 97.77 96.88 73.55 99.55 95.56 95.11 77.05 97.78 75.56 73.87 73.09 75.11
Overall PBR + Train-Def 93.75 85.41 45.09 96.75 72.60 77.53 43.65 92.91 91.11 85.03 53.00 92.59
PBR + Train-Def w/ Aug 89.78 84.37 54.25 96.36 79.20 82.45 56.63 94.73 91.06 90.08 71.17 92.00
PBR + Train-Var 95.26 94.33 73.56 97.44 94.53 94.23 78.40 97.82 91.06 90.50 86.65 91.07
In Table 4, AE consistently exhibits lower performance on Test-Var compared to Test-Def (captured at the same 50% brightness level as Train-Def), regardless of the model or training strategy. This clearly illustrates the necessity of evaluating model robustness under realistic illumination shifts. Applying digital augmentation during training generally improves AE performance on Test-Var, yet adversely affects certain cases (e.g., ZebraPose for Pringles and GDRNPP for Tincase), suggesting that augmentation alone may not reliably enhance robustness. In contrast, training models on physically collected data (Train-Var, which is 3.5 times larger than Train-Def) consistently yields greater performance gains compared to digital augmentation, highlighting the inherent limitations of synthetic simulation in accurately replicating real-world environmental variations. Given that physically collecting large-scale, diverse datasets is often challenging in practice, this performance gap underscores the need for an efficient test-time adaptation methods to effectively address sensor and environmental shifts.
Model-friendly sensor control vs. Traditional auto-exposure. Within Test-Var, AE significantly outperforms Rand, showing the critical role of intentional sensor control in managing environmental variability. Furthermore, Oracle consistently surpasses AE across all models and training strategies, verifying that the auto-exposure mechanism – typically optimized for human visual perception – is not necessarily optimal for model performance. Thus, substantial potential exists for model-specific sensor adaptation at test time, without modifying model architectures or training schemes. Notably, Oracle achieves over 13.4% average performance improvement compared to AE for RGB-based models (ZebraPose and GDRNPP).
Test-time sensor control vs. Traditional generalization approaches. Oracle without augmentation (i.e., the first training scheme) consistently surpasses AE with augmentation, and even outperforms AE trained on Train-Var in most scenarios. In the worst case (GDRNPP), Oracle trained on Train-Def remains comparable to AE trained on Train-Var, showing less than a 2% accuracy gap. These superior results, relative to traditional generalization strategies, such as augmentation or manual scaling training data, demonstrate the substantial potential of test-time sensor control for 6D pose estimation in dynamic real-world environments. Consequently, mitigating covariate shifts during inference through sensor optimization may be more effective, or complementary, compared to exclusively focusing on model generalization to all possible covariate variations, as currently pursued by foundational-model approaches.
4.3 Analysis on Multimodal RGB-D Sensor Control
To comprehensively analyze the impact of multimodal sensor control under environmental variations, we train HiPose (an RGB-D model) following the first training scheme (PBR + Train-Def, captured at 50% brightness level) and evaluate it on the Test-Var split. To specifically assess high-precision pose estimation, we report AUC@[0:0.1], which measures the area under the ADD recall curve within 10% of the object diameter. We consider five conditions: (1) Baseline utilizes the default settings for both RGB and depth sensors; (2) Depth-Only applies Oracle selection for the depth sensor, while maintaining auto-exposure settings for RGB; (3) RGB-Only applies Oracle selection for the RGB sensor, while using the Default mode for the depth sensor; (4) Oracle-Fixed selects the single best-performing fixed RGB-D configuration and applies it across all test scenes; (5) Oracle-Dynamic applies Oracle selection to both RGB and depth sensors for each scene.
@@ 4.3 Analysis on Multimodal RGB-D Sensor Control
Table 5: AUC performance of various sensor control methods for HiPose, trained on PBR + Train-Def and tested on Test-Var.
Object Baseline RGB: Auto Depth: Default Depth-Only RGB: Auto Depth: Oracle RGB-Only RGB: Oracle Depth: Default Oracle-Fixed Best Fixed Param. Oracle-Dynamic RGB: Oracle Depth: Oracle
Spray 86.80 88.04(+1.24) 88.87(+2.07) 86.80 89.92(+3.12)
Pringles 73.65 76.10(+2.45) 78.92(+5.27) 74.68(+1.03) 81.42(+7.77)
Tincase 56.07 58.63(+2.56) 63.60(+7.53) 56.10(+0.03) 66.92 (+10.85)
Overall 72.17 74.25(+2.08) 77.13(+4.95) 72.53(+0.53) 79.42(+7.24)
Table 5 shows that all four sensor-control methods outperform Baseline across every scenario. Among these methods, Oracle-Fixed yields the smallest improvement, offering only marginal gains over Baseline despite controlling multimodal sensor parameters. This result highlights the impracticality of relying on a single, fixed sensor configuration to handle diverse environmental variations; instead, sensor parameters must be adapted dynamically on a per-scene basis.
Complementary role of RGB and Depth. Interestingly, Depth-Only achieves a notable AUC improvement of 2.08% compared to Baseline, demonstrating consistent benefits across all objects. Even though the model was trained exclusively with the Default depth-capture mode – suggesting that employing alternative modes at inference might introduce covariate shifts – allowing flexible depth-mode selection still substantially enhances pose estimation performance. RGB-Only further improves performance relative to Depth-Only, particularly for symmetric objects (Pringles and Tincase). This suggests that RGB surface information may provide crucial cues for accurately estimating the poses of symmetric objects.
Multimodal joint control vs. Unimodal control. Oracle-Dynamic achieves the best overall performance, improving AUC by 7.24% over Baseline and by 2.95% over the second-best RGB-Only method. Notably, Oracle-Dynamic exhibits the most significant gain for Tincase (10.85% above Baseline), the most challenging object in our dataset. These results indicate that RGB and depth modalities provide complementary information, and that jointly adapting both sensors creates synergy exceeding that achievable by independently adapting individual modalities. Thus, our findings highlight the importance of deeper investigations into multimodal sensor adaptation strategies and understanding which combinations of RGB-D inputs are most beneficial for model performance.
4.4 Qualitative Analysis
Analysis on RGB sensor control. Figure 5 illustrates two representative test scenes for GDRNPP, involving the Tincase and Pringles objects, respectively. In both cases, AE results in incorrect pose predictions (via GDRNPP), while Oracle (RGB-only) successfully estimates accurate object poses. These scenes were captured at the lowest brightness level, representing the most challenging environmental condition. These examples highlight that optimal RGB sensor parameters, especially gain values in these cases, vary depending on both object type and object pose, underscoring the necessity of per-scene test-time sensor control. Furthermore, the RGB images selected by Oracle and AE appear visually indistinguishable from a human perspective. This indicates that the optimal direction for model-friendly sensor control is not always intuitively aligned with human visual perception, emphasizing the importance of sensor control strategies specifically tailored to model performance rather than subjective human assessment.
Analysis on depth sensor control. Figure 6 illustrates two representative test scenes evaluated with HiPose, involving the Tincase and Pringles objects, respectively. Both scenes were captured at the lowest brightness level, representing the most challenging environmental conditions. While the RGB sensor parameters were fixed, we compared pose estimation performance across different depth-capture modes. In both cases, the Default depth mode resulted in incorrect pose predictions, whereas Oracle (Depth-only) successfully produced accurate object poses. These examples clearly demonstrate that different depth-capture modes yield varying depth-map qualities, affecting pose estimation outcomes. Furthermore, the optimal depth-capture mode depends on the specific object type and scene configuration, underscoring the importance of adaptive, scene-dependent depth-sensor control strategies.
Figure 5: RGB visualization of pose estimation results under auto exposure and sensor control settings. The red line represents the ground truth pose, and the green line indicates the predicted pose.
Figure 6: Visualization of pose estimation results showing that depth modes other than the default yield better performance under a fixed RGB configuration (E2500, G16). Red lines denote ground truth poses, while green lines represent predicted poses.

### v2

@@ 4 Experiments
Models. To quantify the effect of illumination and sensor variations—and the gains from test‑time sensor control—relative to conventional training–inference paradigms (large‑scale pretraining and standard supervised pipelines), we evaluate several state‑of‑the‑art methods:
  • Pretrained, generalizable models (unseen objects; RGB-D): GigaPose (Nguyen et al., 2024), SAM-6D (Lin et al., 2024a), and FoundationPose (Wen et al., 2024). GigaPose and SAM-6D were pretrained on the large-scale MegaPose (Labbé et al., 2022), which contains objects from ShapeNet (Chang et al., 2015) and GSO (Downs et al., 2022). FoundationPose uses a custom data-generation pipeline with 3D assets from GSO and Objaverse (Deitke et al., 2023).
  • Instance-level models (known objects): ZebraPose (Su et al., 2022), an RGB-based method with a ResNet-34 backbone; GDRNPP (Liu et al., 2025), which builds on GDR-Net (Wang et al., 2021) using a ConvNeXt (Liu et al., 2022) backbone and depth-based refinement module; and HiPose (Lin et al., 2024b), an RGB-D approach that fuses ConvNeXt-based RGB features with RandLA-Net (Hu et al., 2020) depth features. These models are trained on SenseShift6D.
Sensor control. We consider five regimes to evaluate the impact of sensor control: (1) Baseline: default settings for both RGB (e,g, auto-exposure) and depth sensors; (2) Depth-Only: Oracle selection for depth presets and default auto-exposure for RGB; (3) RGB-Only: Oracle selection for RGB exposure and gain, with the Default mode for the depth sensor; (4) Oracle-Fixed: a single best-performing fixed RGB-D configuration, applied to all test scenes; (5) Oracle-Dynamic: per-scene Oracle selection over both RGB and depth. Here, Oracle represents an upper bound obtained by evaluating the full sensor-configuration grid and counting a scene as correct if any configuration yields a correct prediction.
Metrics. We use the standard Average Distance (ADD) metric (Hinterstoisser et al., 2012b), which measures the average point-wise distance between model points transformed by the predicted pose and the ground-truth pose. Based on this metric, we report two evaluation measures: AUC@[0:0.1], which corresponds to the area under the ADD recall curve up to a threshold of 10% of the object diameter, and AR@5, which denotes the recall at a threshold of 5% of the object diameter. Due to space limits, additional metrics—average recall of VSD, MSSD and MSPD (Hodan et al., 2018a)—are reported in Appendix D.
4.1 Analysis on Pretrained, Generalizable Pose Estimation Models
Generalizable 6D pose estimators are designed to cope with clutter, occlusion, and novel objects by leveraging large‑scale pretraining. SenseShift6D removes these scene‑level confounders—single object, clean background, no occlusion—which isolates the effect of environmental and sensor shifts. Evaluating GigaPose, SAM‑6D, and FoundationPose on SenseShift6D thus tests whether robustness acquired through pretraining transfers to an orthogonal and underexplored axis: variations in illumination and RGB/depth sensor parameters that are largely absent from existing benchmarks.
Table 4 presents the impact of illumination and sensor variations on 6D pose estimation, evaluated with ground-truth object segmentations. Given that GigaPose and FoundationPose assume ground-truth masks, while SAM-6D includes its own image segmentation model (ISM), we also provide the results of SAM-6D’s full pipeline in Appendix C for completeness.
Table 4: AUC@[0:0.1] performance of various sensor control methods for unseen object pose estimation models: GigaPose, SAM-6D, FoundationPose.
Model Object
Baseline
RGB: Auto
Depth: Default
Depth-Only
RGB: Auto
Depth: Oracle
RGB-Only
RGB: Oracle
Depth: Default
Oracle-Fixed
Best Fixed Param.
Oracle-Dynamic
RGB: Oracle
Depth: Oracle
GigaPose
 (Nguyen et al., 2024)
Spray 63.94 72.41 (+8.47) 76.42 (+12.48) 67.98 (+4.04) 81.24 (+17.30)
Pringles 20.91 46.74 (+25.83) 46.13 (+25.22) 30.82 (+9.91) 70.54 (+49.63)
Tincase 64.49 72.67 (+8.18) 69.32 (+4.83) 65.51 (+1.02) 76.41 (+11.92)
Sandwich 64.94 71.92 (+6.98) 76.09 (+11.15) 67.90 (+2.96) 83.88 (+18.94)
Mouse 56.62 67.40 (+10.78) 67.02 (+10.40) 53.46 (-3.16) 76.71 (+20.09)
Overall 54.18 66.23 (+12.05) 67.00 (+12.82) 57.13 (+2.95) 77.76 (+23.58)
SAM-6D
 (Lin et al., 2024a)
Spray 79.49 89.41 (+9.92) 87.73 (+8.24) 81.22 (+1.73) 90.80 (+11.31)
Pringles 63.48 79.17 (+15.69) 80.86 (+17.38) 64.54 (+1.06) 83.98 (+20.50)
Tincase 72.26 83.02 (+10.76) 85.49 (+13.23) 68.85 (-3.41) 86.53 (+14.27)
Sandwich 78.39 80.85 (+2.46) 81.59 (+3.20) 78.38 (-0.01) 83.03 (+4.64)
Mouse 53.82 68.24 (+14.42) 74.63 (+20.81) 57.56 (+3.74) 77.61 (+23.79)
Overall 69.49 80.14 (+10.65) 82.06 (+12.57) 70.11 (+0.62) 84.39 (+14.90)
FoundationPose
 (Wen et al., 2024)
Spray 84.63 86.52 (+1.89) 87.60 (+2.97) 83.96 (-0.67) 89.01 (+4.38)
Pringles 34.26 50.72 (+16.46) 50.78 (+16.52) 37.73 (+3.47) 74.69 (+40.43)
Tincase 39.55 58.41 (+18.86) 70.03 (+30.48) 38.25 (-1.30) 80.95 (+41.40)
Sandwich 83.17 86.40 (+3.23) 85.12 (+1.95) 83.14 (-0.03) 88.11 (+4.94)
Mouse 70.22 74.34 (+4.12) 75.86 (+5.64) 70.69 (+0.47) 79.98 (+9.76)
Overall 62.37 71.28 (+8.91) 73.88 (+11.51) 62.76 (+0.39) 82.55 (+20.18)
Figure 5: Overall standard deviation of AUC computed across all brightness and scenes evaluated under Baseline, Oracle-Fixed, Oracle-Dynamic.
Default, Fixed, and Adaptive Multimodal Sensor Configurations. Baseline (auto‑exposure RGB + Default depth) yields the lowest accuracy among the five regimes. Controlling a single modality already helps: RGB‑Only improves by +12.3 pp or more across all models, indicating that factory auto‑exposure—tuned for human perception—is not model‑optimal. Depth‑Only also improves performance by +10.5 pp or more, showing that the default depth preset is likewise suboptimal. Furthermore, Oracle-Dynamic (per-scene RGB-and-depth selection) achieves the best results, with +19.5 pp over Baseline and +7.2 pp over the second-best RGB-Only, demonstrating complementary cues and synergy when both sensors are adapted jointly. In contrast, Oracle-Fixed (a single global configuration) provides only +1.3 pp on average—and can even underperform Baseline (e.g., SAM-6D on Tincase)—underscoring that adaptive, per-scene control is required. Crucially, all gains are obtained without any model retraining, highlighting the practical potential of model-specific sensor adaptation at test time. These results motivate deeper study of multimodal test-time sensor control and of which RGB-D combinations most benefit each model.
Impact on Generalization and Stability. Under Baseline, we observe substantial object‑wise disparity: the average gap between the best‑ and worst‑performing objects is 40.0 pp. Such gaps arise when objects lack distinctive cues for pose disambiguation or deviate geometrically from pretraining distributions, indicating that current generalizable methods still struggle on some novel objects. In contrast, Oracle-Dynamic markedly reduces this disparity to 13.6 pp. Moreover, Figure 5 shows that per‑scene sensor control reduces scene‑level variability under the same object. The overall AUC standard deviation decreases by 8.8 pp relative to Baseline, with GigaPose on the Sandwich showing the largest reduction of 24.2 pp. This indicates higher accuracy and greater stability across diverse scenes.
Together, these findings demonstrate that adapting inputs via sensor control enhances robustness precisely where large‑scale pretrained models tend to fail. While prior advances have emphasized clutter, occlusion, and category novelty, our results reveal that environmental and sensor-setting variation constitute a distinct—and underexplored—axis of difficulty warranting dedicated investigation. Qualitative examples of pretrained, generalizable models are presented in Appendix E.1.
4.2 Analysis on Instance-Level Pose Estimation Models
Instance‑level methods follow the standard train–test setup in which the training and test sets share the same object domain. Building on Section 4.1, our goal here is twofold: (i) to test whether environmental and sensor variations remain challenging once the unseen‑object factor is removed, and (ii) to assess whether test‑time sensor control remains effective even when the training data already includes various augmentations or explicit physical sensor and illumination variation.
@@ 4.2 Analysis on Instance-Level Pose Estimation Models
For training, we generated 10k photorealistic (PBR) synthetic images per object using BlenderProc (Denninger et al., 2023). Each model was trained independently on a per-object basis under three schemes: (1) PBR + Train-Def (default physical images only), (2) PBR + Train-Def with augmentation (including Gaussian blur, linear contrast adjustment, brightness enhancement and others; see Appendix A), and (3) PBR + Train-Var (full sensor and lighting variation). In all experiments, we excluded refinement modules for ZebraPose and GDRNPP, and used the full pipeline for HiPose. Since HiPose is the only RGB-D method (ZebraPose and GDRNPP are RGB-only), we report results in two parts. Table 5 evaluates the impact of illumination and RGB sensor variations (with the default depth preset) under three sensor-control settings: Auto (auto-exposure), Rand (average performance across randomly selected exposure-gain pairs), and Oracle (per-scene upper bound over the full RGB grid). Table 6 evaluates HiPose (training scheme 1) on Test-Var under the multimodal regimes introduced in Section 4.1.
@@
Table 5: AR@5 performance of various sensor control methods under different training and test settings of SenseShift6D for instance-level pose estimation models: ZebraPose, GDRNPP, HiPose.
ZebraPose (Su et al., 2022) GDRNPP (Liu et al., 2025) HiPose (Lin et al., 2024b)
@@
Object Train Auto Auto Rand Oracle Auto Auto Rand Oracle Auto Auto Rand Oracle
Spray PBR + Train-Def 100.0 96.09 52.26 99.51 100.0 96.59 52.02 99.02 100.0 98.73 59.98 100.0
@@
PBR + Train-Var 90.47 89.04 68.35 95.23 90.48 90.00 72.22 97.14 97.62 97.62 86.86 98.10
Tincase PBR + Train-Def 95.56 79.56 52.53 96.44 91.11 81.33 46.58 92.89 97.78 81.51 53.62 96.44
PBR + Train-Def w/ Aug 97.78 89.78 65.46 100.0 80.00 82.22 52.89 95.11 97.78 95.64 86.47 97.78
PBR + Train-Var 98.22 97.60 79.59 100.0 100.0 99.56 82.02 100.0 97.78 97.69 97.76 97.78
Sandwich PBR + Train-Def 94.87 89.23 48.26 100.0 97.44 88.21 47.03 96.41 94.87 87.69 57.21 94.87
PBR + Train-Def w/ Aug 92.31 88.72 57.82 98.97 94.87 92.31 63.72 98.97 94.87 93.74 84.38 94.87
PBR + Train-Var 89.74 89.23 67.15 96.93 97.44 97.44 82.64 100.0 94.87 94.87 94.87 94.87
Mouse PBR + Train-Def 70.73 61.46 27.37 86.34 95.12 76.10 31.44 88.78 90.24 62.93 43.73 87.81
PBR + Train-Def w/ Aug 48.78 49.76 32.56 88.29 78.05 72.68 43.29 86.83 90.24 89.07 82.24 90.24
PBR + Train-Var 46.34 46.34 34.98 73.17 80.49 80.49 66.39 91.22 90.24 89.95 87.17 90.73
Overall PBR + Train-Def 89.37 80.41 43.86 95.41 90.54 80.64 42.54 93.23 96.56 85.37 54.66 95.82
PBR + Train-Def w/ Aug 82.09 76.45 50.46 94.74 86.77 84.96 56.19 94.09 96.10 95.25 83.10 96.58
PBR + Train-Var 84.47 83.86 65.77 92.58 93.19 93.01 77.84 97.38 96.10 96.03 93.33 96.30
Sensor and Environmental Shift. In Table 5, when trained with Train-Def, Auto performance is consistently lower on Test-Var than Test-Def, reflecting sensitivity to illumination shifts. In contrast, Oracle on Test‑Var not only surpasses Auto on Test‑Var but often exceeds Auto on Test‑Def, showing that adapting RGB sensor settings at test time both mitigates domain shift and unlocks higher accuracy. Table 6 (RGB-D) reveals that while all sensor‑control methods improve over Baseline, depth‑only control provides marginal gains relative to RGB control—unlike the trend in Table 4. This is expected because: (i) instance‑level training for HiPose exposes the model to the exact objects, yielding object‑specific optimization that leaves less head‑room for improved depth measurements; (ii) training already uses depth maps from the same domain as Baseline; and (iii) depth sensing is inherently less sensitive to illumination changes than RGB. Even with identical objects and backgrounds, however, performance still varies under environmental and sensor changes.
@@
Table 6: AUC@[0:0.1] performance of various sensor control methods for HiPose, trained on PBR + Train-Def and tested on Test-Var.
Object
Baseline
RGB: Auto
Depth: Default
Depth-Only
RGB: Auto
Depth: Oracle
RGB-Only
RGB: Oracle
Depth: Default
Oracle-Fixed
Best Fixed Param.
Oracle-Dynamic
RGB: Oracle
Depth: Oracle
Spray 86.92 88.17 (+1.25) 88.79 (+1.87) 87.54 (+0.62) 89.87 (+2.95)
Pringles 73.40 76.18 (+2.78) 79.08 (+5.68) 77.02 (+3.62) 81.55 (+8.15)
Tincase 74.78 77.88 (+3.10) 85.30 (+10.52) 83.57 (+8.79) 86.84 (+12.06)
Sandwich 78.68 81.72 (+3.04) 84.06 (+5.38) 82.90 (+4.22) 86.43 (+7.75)
Mouse 63.92 66.41 (+2.49) 72.19 (+8.27) 69.52 (+5.60) 75.21 (+11.29)
Overall 75.54 78.07 (+2.53) 81.88 (+6.34) 80.11 (+4.57) 83.98 (+8.44)
[…]
These results demonstrate the substantial potential of test‑time sensor control relative to conventional generalization strategies (augmentation or scaling training data), even for known objects. Mitigating covariate shift at inference via sensor optimization can be more effective—or complementary—than exclusively attempting to generalize the model to all possible shifts. Qualitative results are provided in Appendix E.2.